# Focusing by Contrastive Attention: Enhancing VLMs’ Visual Reasoning

### v1

@@ 2 Related Work
3 Attention May Go Myopic: When, Why, and How
3.1 Discovery: When Visual Focus Fails
@@
3.2 Root Cause: Why VLMs Go Myopic
@@
3.3 Cascade: How Performance Goes Astray
@@ 5.2 Results
To investigate layer selection effects on attention pattern extraction, we conduct systematic experiments as shown in Table 2. Across all tested model architectures, the layer-wise performance demonstrates the following general ordering: $20,25]\succ[15,20]\succ\text{single layer }25\approx\text{single layer }20\succ\text{single layer }14\gg[10,15]$. This pattern is exemplified by Llava1.5-7B’s performance on TextVQA, where the multi-layer [20,25] achieves a 21.76% improvement, the [15,20] reaches a 17.99% improvement, while the early-layer [10,15] attains only a 2.93% improvement. Multi-layer fusion outperforms single-layer alternatives by capturing complementary information and providing robustness against individual layer randomness. This phenomenon aligns with our findings in Figure 5(b): early layers perform global scanning with high entropy, while middle-to-deep layers focus on task-relevant patterns.
@@ Appendix B Implementation Details
We conduct our experiments on a server with 4 $\times$ NVIDIA RTX A6000 GPUs. In practical implementation, CARVE requires three inference passes; however, the first two passes (extracting general instruction and task-specific question attention maps) can be terminated early. Specifically, when we require attention maps only from layers $\mathcal{L}=[L_{start},L_{end}]$, the first two inference processes can halt upon completing layer $L_{end}$ computation, eliminating the need for full $L_{\text{total}}$ layer forward propagation. The third inference must run to completion to generate the final answer.
@@ C.1 Mathematical Basis of Attention Decomposition
Theorem B.1 (Existence of Attention Decomposition): For any attention distribution $A_{l,t}^{(Q)}(\mathcal{I})\in\mathbb{R}^{N_{v}}_{+}$, there exists a unique decomposition:
@@ C.2 Convexity Analysis of the Optimization Problem
Theorem B.2 (Strict Convexity of Objective Function): The optimization objective
@@ C.3 Derivation and Uniqueness of Closed-form Solution
Theorem B.3 (Closed-form Expression of Optimal Solution): The optimization problem admits a unique global optimum:
@@
By Theorem B.2’s strict convexity, this solution represents the unique global optimum. $\square$
@@ C.4 Error Bounds and Convergence Analysis
Theorem B.4 (Approximation Error Bound): Let $\mathcal{F}_{\text{sem}}(G,\mathcal{I})=\mathbf{1}_{N_{v}}+\epsilon$ where $\|\epsilon\|_{\infty}\leq\delta$. Then the estimation error satisfies:
@@ C.5 Theoretical Selection of Regularization Parameter
Proposition B.5 (Optimal Regularization Parameter): In the mean squared error sense, the optimal regularization parameter satisfies:
@@
Analysis: Defining the bias-variance decomposition:
@@
where the bias term $\text{Bias}(\lambda)=\mathbb{E}[\hat{A}(\lambda)]-\mathcal{F}_{\text{sem}}(Q,\mathcal{I})$ and the variance term $\text{Variance}(\lambda)=\mathbb{E}[(\hat{A}(\lambda)-\mathbb{E}[\hat{A}(\lambda)])^{2}]$.
As $\lambda\rightarrow 0$, variance increases due to numerical instability while bias decreases. Conversely, as $\lambda\rightarrow\infty$, variance decreases but bias increases due to over-smoothing. The optimal $\lambda^{*}$ achieves the bias-variance trade-off point.
@@ C.6 Hierarchical Evolution of Attention Entropy
Theorem B.6 (Monotonicity of Entropy): For a layer sequence $l_{1}<l_{2}<...<l_{n}$, attention entropy satisfies:
@@
Proof sketch: Applying the Data Processing Inequality, we treat each layer as an information processing channel. Since deeper networks progressively extract high-level features and focus on task-relevant regions, information entropy decreases monotonically. This aligns with the principle of maximum entropy: systems tend toward maximum entropy states under constraints, where deeper layers impose stronger task constraints. $\square$

### v2

@@ 2 Related Work
3 Failure to Focus: Phenomenon, Mechanism, and Consequence
3.1 Phenomenon: Under What Conditions Visual Focus Fails
@@
3.2 Mechanism: The Effect of Visual Complexity on Attention Entropy
@@
3.3 Consequence: How Dispersed Attention Impairs Performance
@@ 5.2 Results
To investigate layer selection effects on attention pattern extraction, we conduct systematic experiments as shown in Table 2. Across all tested model architectures, the layer-wise performance demonstrates the following general ordering from best to worst: [20,25], [15,20], single layer 25, single layer 20, single layer 14, and [10,15]. This pattern is exemplified by Llava1.5-7B’s performance on TextVQA, where the multi-layer [20,25] achieves a 21.76% improvement, the [15,20] reaches a 17.99% improvement, while the early-layer [10,15] attains only a 2.93% improvement. Multi-layer fusion outperforms single-layer alternatives by capturing complementary information and providing robustness against individual layer randomness. This phenomenon aligns with our findings in Figure 5(b): early layers perform global scanning with high entropy, while middle-to-deep layers focus on task-relevant patterns.
@@ A.2 Explanation
  • Visual Complexity ($\mathcal{T}_{c}(\mathcal{I})$, $\mathcal{C}_{c}(\mathcal{I})$): Visual complexity quantifies the inherent characteristics of an image that can interfere with VLMs’ attention mechanisms, decomposed into two orthogonal dimensions. Texture complexity $\mathcal{T}_{c}(\mathcal{I})\in[0,1]$ measures the density of edge information using Canny edge detection, where higher values indicate more intricate patterns, object boundaries, and structural details. Color complexity $\mathcal{C}_{c}(\mathcal{I})\in[0,1]$ captures the diversity of hue distribution in HSV color space through Shannon entropy, where higher values reflect greater chromatic variation.
@@ Appendix B Implementation Details
We conduct our experiments on a server with 4 $\times$ NVIDIA RTX A6000 GPUs. $\tau$ is set to 0.05. In practical implementation, CARVE requires three inference passes; however, the first two passes (extracting general instruction and task-specific question) can be terminated early. Specifically, when we require attention maps only from layers $\mathcal{L}=[L_{start},L_{end}]$, the first two inference processes can halt upon completing layer $L_{end}$ computation, eliminating the need for full $L_{\text{total}}$ layer forward propagation. The third inference must run completely to generate the final answer.
@@ C.1 Mathematical Basis of Attention Decomposition
Theorem C.1 (Existence of Attention Decomposition): For any attention distribution $A_{l,t}^{(Q)}(\mathcal{I})\in\mathbb{R}^{N_{v}}_{+}$, there exists a unique decomposition:
@@ C.2 Convexity Analysis of the Optimization Problem
Theorem C.2 (Strict Convexity of Objective Function): The optimization objective
@@ C.3 Derivation and Uniqueness of Closed-form Solution
Theorem C.3 (Closed-form Expression of Optimal Solution): The optimization problem admits a unique global optimum:
@@
By Theorem C.2’s strict convexity, this solution represents the unique global optimum. $\square$
@@ C.4 Error Bounds and Convergence Analysis
Theorem C.4 (Approximation Error Bound): Let $\mathcal{F}_{\text{sem}}(G,\mathcal{I})=\mathbf{1}_{N_{v}}+\epsilon$ where $\|\epsilon\|_{\infty}\leq\delta$. Then the estimation error satisfies:
@@ C.5 Theoretical Selection of Regularization Parameter
Proposition C.5 (Optimal Regularization Parameter): The optimal regularization parameter $\lambda^{*}$ that minimizes the expected mean squared error satisfies:
@@
Proof: From Theorem C.3, the estimator takes the form:
$\hat{A}_{i}(\lambda)=\frac{A^{(Q)}_{i}}{A^{(G)}_{i}+\lambda}=\frac{\mathcal{F}_{\text{vis},i}\cdot\mathcal{F}_{\text{sem},i}}{\mathcal{F}_{\text{vis},i}+\lambda}$ (C.12)
The mean squared error decomposes as:
@@
where $\text{Bias}(\lambda)=\mathbb{E}[\hat{A}(\lambda)]-\mathcal{F}_{\text{sem}}(Q,\mathcal{I})$ and $\text{Variance}(\lambda)=\mathbb{E}[(\hat{A}(\lambda)-\mathbb{E}[\hat{A}(\lambda)])^{2}]$.
For the bias term, assuming $\mathbb{E}[\mathcal{F}_{\text{vis},i}]=\mu_{i}$:
$\text{Bias}_{i}(\lambda)=\mathbb{E}\left[\frac{\mathcal{F}_{\text{vis},i}\cdot\mathcal{F}_{\text{sem},i}}{\mathcal{F}_{\text{vis},i}+\lambda}\right]-\mathcal{F}_{\text{sem},i}\approx-\frac{\lambda\cdot\mathcal{F}_{\text{sem},i}}{\mu_{i}+\lambda}$ (C.14)
Thus $|\text{Bias}_{i}(\lambda)|=O(\lambda)$ as $\lambda\to 0$.
For the variance term, let $\mathcal{F}_{\text{vis},i}=\mu_{i}+\epsilon_{i}$ with $\text{Var}(\epsilon_{i})=\sigma_{i}^{2}$. Taylor expansion yields:
$\text{Var}(\hat{A}_{i}(\lambda))\approx\frac{\mathcal{F}_{\text{sem},i}^{2}\mu_{i}^{2}\sigma_{i}^{2}}{(\mu_{i}+\lambda)^{4}}$ (C.15)
Therefore $\text{Var}(\hat{A}_{i}(\lambda))=O(1/\lambda^{2})$ as $\lambda\to 0$.
The component-wise MSE becomes:
$\text{MSE}_{i}(\lambda)=\frac{\lambda^{2}\cdot\mathcal{F}_{\text{sem},i}^{2}}{(\mu_{i}+\lambda)^{2}}+\frac{\mathcal{F}_{\text{sem},i}^{2}\mu_{i}^{2}\sigma_{i}^{2}}{(\mu_{i}+\lambda)^{4}}$ (C.16)
Setting $\frac{d\text{MSE}_{i}}{d\lambda}=0$ and solving yields:
$\lambda_{i}^{*}=\mu_{i}\left(\sqrt{1+2\sigma_{i}^{2}/\mu_{i}^{2}}-1\right)\approx\frac{\sigma_{i}^{2}}{\mu_{i}}$ (C.17)
for small noise-to-signal ratio. The global optimum requires minimizing $\sum_{i=1}^{N_{v}}\text{MSE}_{i}(\lambda)$. $\square$
Corollary C.5.1 (Numerical Stability): For any $\lambda>0$, the condition number of the regularized problem satisfies:
$\kappa(\lambda)=\frac{\max_{i}(A^{(G)}_{i}+\lambda)}{\min_{i}(A^{(G)}_{i}+\lambda)}\leq\frac{\max_{i}A^{(G)}_{i}+\lambda}{\lambda}$ (C.18)
Proof: The bound follows directly from the definition of condition number and the positivity of $A^{(G)}_{i}$ and $\lambda$. The regularization ensures $\kappa(\lambda)<\infty$, guaranteeing numerical stability. $\square$
Remark: The regularization parameter $\lambda$ serves dual purposes: controlling the bias-variance trade-off and ensuring numerical stability. As $\lambda\to 0$, the estimator becomes unbiased but exhibits high variance and potential numerical instability when $A^{(G)}_{i}\approx 0$. Conversely, as $\lambda\to\infty$, the estimator becomes increasingly biased toward zero but achieves maximum stability. The optimal choice $\lambda^{*}\propto\sigma^{2}/\mu$ balances these competing objectives, where $\sigma^{2}$ represents the noise variance and $\mu$ the signal mean. In practice, cross-validation on a held-out set provides robust estimation of $\lambda^{*}$.
@@ C.6 Hierarchical Evolution of Attention Entropy
Theorem C.6 (Monotonicity of Entropy): For a layer sequence $l_{1}<l_{2}<...<l_{n}$, attention entropy satisfies:
@@
Proof: Applying the Data Processing Inequality, we treat each layer as an information processing channel. Since deeper networks progressively extract high-level features and focus on task-relevant regions, information entropy decreases monotonically. This aligns with the principle of maximum entropy: systems tend toward maximum entropy states under constraints, where deeper layers impose stronger task constraints. $\square$